OpenStudy (anonymous):

Translate the problem situation to a system of equations. Do not solve. The difference between two numbers is 1. Three times the smaller number plus the larger number is 29. What are the numbers? A. x - y = 1; x + 3z = 29 B. x - y = 1; 3x + y = 29 C. x - y = 1; x + 3y = 29 D. x + y = 1; x + 3y = 29

OpenStudy (phoenixfire):

What do you think the first one is? "Difference between two numbers is 1"

OpenStudy (anonymous):

its B

OpenStudy (phoenixfire):

@terminallychill It's not B... And don't give out the answer.

OpenStudy (anonymous):

its A

OpenStudy (phoenixfire):

@mollyann It's not A either.

OpenStudy (phoenixfire):

Answer my question. Difference between two numbers is 1. What's the equation?

OpenStudy (anonymous):

definitely D

OpenStudy (anonymous):

yea its D

OpenStudy (phoenixfire):

@terminallychill It's not D. Not I've given the answer.. It's C

OpenStudy (anonymous):

\m/

OpenStudy (phoenixfire):

Difference between two number is 1: x-y=1... The smaller number is y. 3 times the smaller number + the larger number is 29.: x+3y=29

OpenStudy (phoenixfire):

y is the smaller number because x - y > 0 if y < x x - y < 0 if y > x

OpenStudy (anonymous):

its not C

OpenStudy (anonymous):

my test said thats wrong

OpenStudy (phoenixfire):

Show me your test... there's only two variables. x and y. from that you know A is wrong as it includes a z. It wants a difference between two numbers, so it's x-y. This rules out D as x+y is not a difference. 3 times the SMALLER number. this is in C as y is LESS THAN x for the difference x-y=1